# Optimization of the Downlink Spectral- and Energy-Efficiency of RIS-aided Multi-user URLLC MIMO Systems

### v1

@@ III-B Optimizing the RIS Elements
Substituting the surrogate functions for the rates, i.e., the $\hat{r}_{lk}$s, in (III-B) yields
@@
which is convex if $\mathcal{T}$ is a convex set, i.e. when $\mathcal{T}_{U}$ is considered. For $\mathcal{T}_{U}$, the proposed scheme achieves convergence to a stationary point of (II-D). Note that the surrogate functions $\hat{f}_{i}\left(\left\{\mathbf{P}^{(t)}\right\}\!,\!\{\bf{\Psi}\}\right)$ are concave in $\{{\bf\Psi}\}$ even if they are linear functions of the EE metrics. The reason is that the powers (transmit covariance matrices) are fixed, and thus, the EE metrics are not fractional functions of $\{{\bf\Psi}\}$.
@@
The constraint $|\psi_{m_{n}}|^{2}\leq 1$ is convex; however, $|\psi_{m_{n}}|^{2}\geq 1$ is not, which makes (III-B) a non-convex problem. Thus, we have to approximate (41) with a convex constraint to make (III-B) convex. To this end, we employ the convex-concave procedure (CCP) and rewrite (41) as
@@
where $\delta>0$. Now, we can approximate (III-B) as
@@
which is convex. We denote the solution of (III-B) as $\{{\bf\Psi}^{(\star)}\}=\{{\bf\Psi}^{(\star)}_{1},{\bf\Psi}^{(\star)}_{2},\cdots,{\bf\Psi}^{(\star)}_{M}\}$. Because of the relaxation in (43), it may happen that $\psi_{m_{n}}^{(\star)}$, i.e., the $n$-th coefficient of the diagonal matrix ${\bf\Psi}^{(\star)}_{m}$, does not satisfy $|\psi_{m_{n}}|=1$. Therefore, we normalize $\{{\bf\Psi}^{(\star)}\}$ as
@@
For $\mathcal{T}=\mathcal{T}_{I}$, our proposed framework converges since a non-decreasing sequence of objective functions (OF) $f_{0}$ is generated. For $\mathcal{T}=\mathcal{T}_{U}$, our framework converges to a stationary point of (II-D) because $\mathcal{T}_{U}$ is convex. We summarize our algorithm for maximizing the minimum EE with $\mathcal{T}_{U}$ in Algorithm I.
@@
           Compute $\{{\bf Q}\}$ by solving (III-A1), i.e., by running
@@
III-B1 STAR-RIS using the MS scheme
The solution of (III-B) for STAR-RIS with the MS scheme is very similar to the solution for a reflective RIS. In this case, we can still use the surrogate functions in Corollary 1 to make the rates a jointly concave function of the STAR-RIS coefficients, i.e., $\{{\bf\Psi}^{t}\}$ and $\{{\bf\Psi}^{r}\}$. For $\mathcal{T}_{SU}$, we can update the STAR-RIS parameters by solving (III-B), and the algorithm obtains a stationary point of (II-D). For $\mathcal{T}_{SI}$, we have to “convexify” the constraint in (8), which can be done by rewriting it as the two convex constraints in [43, Eq. (34)] and [43, Eq. (36)]. Then we can update the STAR-RIS coefficients similar to the proposed scheme for the reflective RIS.
@@ III-C Computational Complexity Analysis
Each iteration of our proposed framework consists of two steps. In the first step, we optimize the transmit covariance matrices by solving (III-A), which is convex when the minimum rates of the users are maximized. We solve the convex problem in (III-A) by numerical optimization tools. To numerically solve a convex optimization problem, the number of Newton iterations increases proportionally with the square root of the number of its constraints [45, Chapter 11], which is equal to $L(K+1)$ in (III-A) for the maximization of the minimum rate. Note that the maximization of the minimum rate can be written as in [11, Eq. (30)] and thus, has $L(K+1)$ constraints, considering the power budget in (32d). Now, we provide an approximate upper bound for the number of multiplications to find a solution in each Newton iteration. To solve each Newton iteration, $LK$ surrogate functions have to be computed for the rates, $\tilde{r}_{lk}$. The surrogate rates $\tilde{r}_{lk}$ in Lemma 5 are quadratic in $\{{\bf Q}\}$, and the computational complexity to compute each surrogate rate can be approximated as $\mathcal{O}\left[LKN_{BS}^{2}(2N_{BS}+N_{u})\right]$. Note that the coefficients in (27) can be computed once at the beginning of the Newton iterations, and there is no need to recompute them in each Newton iteration to reduce the overall computational complexity of the framework. Finally, the computational complexity to update $\{{\bf P}\}$ can be approximated as $\mathcal{O}\left[L^{2}K^{2}N_{BS}^{2}\sqrt{L(K+1)}(2N_{BS}+N_{u})\right]$.
@@
Now, we derive an approximation for the number of the multiplications needed to update $\{{\bf\Psi}\}$. To this end, we have to calculate the number of multiplications for solving the surrogate optimization problem in (III-B), which is convex for the feasibility set $\mathcal{T}_{U}$. The number of constraints in (III-B) is equal to $LK+MN_{RIS}$. Thus, the number of the Newton iterations grows with $\sqrt{LK+MN_{RIS}}$. To solve each Newton iteration, we have to compute $LK$ surrogate functions for the rates, $\hat{r}_{lk}$, as well as $L^{2}K$ equivalent channels, according to (2). To compute each channel, ${\bf H}_{lk,i}$, $\forall l,k,i$, there are approximately $MN_{u}N_{BS}N_{RIS}$ multiplications, since the matrices ${\bf\Psi}_{m}$ are diagonal, which reduces the computational complexity. Moreover, the structure of the rates in (38) is very similar to the rates in (27). Hence, the computational complexity of calculating $\hat{r}_{lk}$ in (38) is on the same order of the computational complexity of attaining $\tilde{r}_{lk}$ as in (27) and can be approximated as $\mathcal{O}\left[LKN_{BS}^{2}(2N_{BS}+N_{u})\right]$. Finally, the computational complexity of updating $\{{\bf\Psi}\}$ can be approximated as $\mathcal{O}[L^{2}K\sqrt{LK+MN_{RIS}}(KN_{BS}^{2}(2N_{BS}+N_{u})+MN_{u}N_{BS}N_{RIS})]$. Assuming that the maximum number of iterations is equal to $T$, the computational complexity of solving the maximization of the minimum rate for $\mathcal{T}_{U}$ using our framework is $T$ times the summation of the computational complexities of updating $\{{\bf P}\}$ and $\{{\bf\Psi}\}$.
@@ IV Numerical results
In this section, we provide numerical results based on Monte Carlo simulations. To this end, we consider a two-cell system with one RIS per cell, similar to [43, Fig. 2]. Moreover, the locations and heights of the users/BSs/RISs are chosen similar to [43]. We assume that the power budgets of the BSs are equal to $P$. To model the small-scale fading of the channels, we assume that the links through the RISs are line of sight (LoS), and therefore, follow Rician fading with Rician factor $3$. On the other hand, we assume that the direct links between the BSs and users are of non-LoS (NLoS) nature and consequently, follow a Rayleigh distribution. The large-scale fading and other simulation parameters are based on [43, 25]. In the following, we provide numerical results for SE and EE maximization in Section IV-A and Section IV-B, respectively. The schemes considered in this section are as follows:
[…]
@@ IV-A1 Impact of power budget
In Fig. 2, we show the average max-min rate and sum rate versus $P$ for $N_{{BS}}=4$, $N_{u}=4$, $K=2$, $L=2$, $M=2$, $n_{t}=256$, $\epsilon=10^{-5}$, and $N_{{RIS}}=20$. In this figure, the RISs significantly improve the min-max rate and the sum rate, when the RIS elements are optimized. Surprisingly, employing RISs with random elements degrades the max-min rate. Moreover, we can observe that the benefits of RIS increase with $P$ in these examples. Additionally, the multi-stream scheme substantially improves the SE in these two examples.
@@ IV-A2 Impact of packet length
Fig. 3 shows the average max-min rate and the relative performance improvement by RIS versus $n_{t}$ for $P=10$ dB, $N_{{BS}}=4$, $N_{u}=3$, $K=2$, $L=2$, $M=2$, $\epsilon=10^{-5}$, and $N_{{RIS}}=20$. Observe that a RIS can substantially increase the average max-min rate. The benefits of deploying RISs decrease with $n_{t}$. Indeed, the shorter the packet length is, the higher the relative improvements provided by RISs can be. As indicated, the packet length correlates with the level of stringency in the latency constraint. Shorter packet lengths are needed, when the latency constraint is more stringent. Thus, this result shows that the RIS benefits increase as the latency constraint becomes more stringent. In other words, RISs can even be more beneficial for URLLC systems. Moreover, the rates increase with $n_{t}$ and converge to the Shannon rate when $n_{t}$ becomes higher. Furthermore, we can observe that employing multi-stream data transmission substantially increases the average max-min rate for all the values of $n_{t}$.
@@ IV-B1 Impact of $P_{c}$
Fig. 8 shows the average max-min EE and GEE versus $P_{c}$ for $N_{{BS}}=5$, $N_{u}=5$, $K=2$, $L=2$, $M=2$, $n_{t}=256$ bits, $\epsilon=10^{-5}$, and $N_{{RIS}}=20$. Note that we use the term max-min EE to refer to the highest minimum EE. As it can be observed, RISs significantly increase the average max-min EE and GEE with FBL. Interestingly, RISs may reduce the EE, if their elements are random. However, our proposed algorithms can statistically enhance the EE of RIS-aided scenarios. For instance, in the particular example of Fig. 8a, a RIS provides more than $10\%$ improvements over the systems disregarding the RISs for all the values of $P_{c}$ considered. Additionally, the multi-stream scheme significantly outperforms the single-stream data transmission in the both examples.

### v2

@@ III-B1 Reflective RISs
Substituting the surrogate functions for the rates, i.e., the $\hat{r}_{lk}$s, in (40) yields
@@
which is convex if $\mathcal{T}$ is a convex set, i.e. when $\mathcal{T}_{U}$ is considered. For $\mathcal{T}_{U}$, the proposed scheme achieves convergence to a stationary point of (24). Note that the surrogate functions $\hat{f}_{i}\left(\left\{\mathbf{P}^{(t)}\right\}\!,\!\{\bf{\Psi}\}\right)$ are concave in $\{{\bf\Psi}\}$ even if they are linear functions of the EE metrics. The reason is that the powers (transmit covariance matrices) are fixed, and thus, the EE metrics are not fractional functions of $\{{\bf\Psi}\}$.
@@
The constraint $|\psi_{m_{n}}|^{2}\leq 1$ is convex; however, $|\psi_{m_{n}}|^{2}\geq 1$ is not, which makes (42) a non-convex problem. Thus, we have to approximate (44) with a convex constraint to make (42) convex. To this end, we employ the convex-concave procedure (CCP) and rewrite (44) as
@@
where $\delta>0$. Now, we can approximate (42) as
@@
which is convex. We denote the solution of (47) as $\{{\bf\Psi}^{(\star)}\}=\{{\bf\Psi}^{(\star)}_{1},{\bf\Psi}^{(\star)}_{2},\cdots,{\bf\Psi}^{(\star)}_{M}\}$. Because of the relaxation in (46), it may happen that $\psi_{m_{n}}^{(\star)}$, i.e., the $n$-th coefficient of the diagonal matrix ${\bf\Psi}^{(\star)}_{m}$, does not satisfy $|\psi_{m_{n}}|=1$. Therefore, we normalize $\{{\bf\Psi}^{(\star)}\}$ as
@@
For $\mathcal{T}=\mathcal{T}_{I}$, our proposed framework converges since a non-decreasing sequence of objective functions (OF) $f_{0}$ is generated. For $\mathcal{T}=\mathcal{T}_{U}$, our framework converges to a stationary point of (24) because $\mathcal{T}_{U}$ is convex. We summarize our algorithm for maximizing the minimum EE with $\mathcal{T}_{U}$ in Algorithm I.
@@
              Compute $\{{\bf Q}\}$ by solving (34), i.e., by running
@@
III-B2 STAR-RIS using the MS scheme
The solution of (40) for STAR-RIS with the MS scheme is very similar to the solution for a reflective RIS. In this case, we can still use the surrogate functions in Corollary 1 to make the rates a jointly concave function of the STAR-RIS coefficients, i.e., $\{{\bf\Psi}^{t}\}$ and $\{{\bf\Psi}^{r}\}$. For $\mathcal{T}_{SU}$, we can update the STAR-RIS parameters by solving (42), and the algorithm obtains a stationary point of (24). For $\mathcal{T}_{SI}$, we have to “convexify” the constraint in (8), which can be done by rewriting it as the two convex constraints in [36, Eq. (34)] and [36, Eq. (36)]. Then we can update the STAR-RIS coefficients similar to the proposed scheme for the reflective RIS.
@@ III-C Computational Complexity Analysis
Each iteration of our proposed framework consists of two steps. In the first step, we optimize the transmit covariance matrices by solving (33), which is convex when the minimum rates of the users are maximized. We solve the convex problem in (33) by numerical optimization tools. To numerically solve a convex optimization problem, the number of Newton iterations increases proportionally with the square root of the number of its constraints [62, Chapter 11], which is equal to $\sum_{l}(K_{l}+1)$ in (33) for the maximization of the minimum rate. Note that the maximization of the minimum rate can be written as in [14, Eq. (30)] and thus, has $\sum_{l}(K_{l}+1)$ constraints, considering the power budget in (33d). Now, we provide an approximate upper bound for the number of multiplications to find a solution in each Newton iteration. To solve each Newton iteration, $\sum_{l}K_{l}$ surrogate functions have to be computed for the rates, $\tilde{r}_{lk}$. The surrogate rates $\tilde{r}_{lk}$ in Lemma 5 are quadratic in $\{{\bf Q}\}$, and the computational complexity to compute each surrogate rate can be approximated as $\mathcal{O}\left[\sum_{l}\sum_{k}N_{BS,l}^{2}(2N_{BS,l}+N_{u,lk})\right]$. Note that the coefficients in (28) can be computed once at the beginning of the Newton iterations, and there is no need to recompute them in each Newton iteration to reduce the overall computational complexity of the framework. Finally, the computational complexity to update $\{{\bf P}\}$ can be approximated as $\mathcal{O}\left[\mathcal{L}\sum_{l}\sum_{k}N_{BS,l}^{2}\sqrt{\sum_{l}(K_{l}+1)}(2N_{BS,l}+N_{u,lk})\right]$, where $\mathcal{L}=\sum_{l}K_{l}$ is the total number of users in the system.
@@
Now, we derive an approximation for the number of the multiplications needed to update $\{{\bf\Psi}\}$. To this end, we have to calculate the number of multiplications for solving the surrogate optimization problem in (42), which is convex for the feasibility set $\mathcal{T}_{U}$. The number of constraints in (42) is equal to $\sum_{l}K_{l}+\sum_{m}N_{RIS,m}$. Thus, the number of the Newton iterations grows with $\sqrt{\sum_{l}K_{l}+\sum_{m}N_{RIS,m}}$. To solve each Newton iteration, we have to compute $\mathcal{L}$ surrogate functions for the rates, $\hat{r}_{lk}$, as well as $L\mathcal{L}$ equivalent channels, according to (2). To compute each channel, ${\bf H}_{lk,i}$, $\forall l,k,i$, there are approximately $\sum_{m}N_{u,lk}N_{BS,i}N_{RIS,m}$ multiplications, since the matrices ${\bf\Psi}_{m}$ are diagonal, which reduces the computational complexity. Moreover, the structure of the rates in (41) is very similar to the rates in (28). Hence, the computational complexity of calculating $\hat{r}_{lk}$ in (41) is on the same order of the computational complexity of attaining $\tilde{r}_{lk}$ as in (28) and can be approximated as $\mathcal{O}\left[\sum_{l}K_{l}N_{BS,l}^{2}(2N_{BS,l}+N_{u,lk})\right]$. Finally, the computational complexity of updating $\{{\bf\Psi}\}$ can be approximated as $\mathcal{O}[\sum_{l}\mathcal{L}\sqrt{\sum_{l}K_{l}+\sum_{m}N_{RIS,m}}(K_{l}N_{BS,l}^{2}(2N_{BS,l}+N_{u,lk})+\sum_{m}N_{u,lk}N_{BS,i}N_{RIS,m})]$. Assuming that the maximum number of iterations is equal to $T$, the computational complexity of solving the maximization of the minimum rate for $\mathcal{T}_{U}$ using our framework is $T$ times the summation of the computational complexities of updating $\{{\bf P}\}$ and $\{{\bf\Psi}\}$.
III-D Discussion on Extending the Framework to Uplink
The framework can also be applied to the SE and EE maximization of the uplink along with FBL coding, since the structure of the rates with respect to the beamforming matrices and channels is very similar to the DL scenario considered in the paper. The detailed solution for the UL scenario is beyond the scope of this work. However, we provide some insights on how the proposed solutions can be modified to maximize the sum rate in UL communications in a multi-cell multiple-access channel (MAC), while the intercell interference is treated as noise. In this case, the sum rate of users associated with BS $l$ is
$r_{l}=\log|{\bf I}+{\bf D}_{l}^{-1}{\bf S}_{l}|-\frac{Q^{-1}(\epsilon)}{n_{t}}\sqrt{2\text{Tr}({\bf I}-{\bf D}_{l}({\bf D}_{l}+{\bf S}_{l})^{-1})},$ (50)
where ${\bf S}_{l}=\sum_{k}{\bf H}_{lk,l}{\bf Q}_{lk}({\bf H}_{lk,l}{\bf Q}_{lk})^{H}$ is the covariance matrix of the signals decoded at BS $l$, and ${\bf D}_{l}=\sigma^{2}{\bf I}+\sum_{\forall ik,i\neq l}{\bf H}_{lk,l}{\bf Q}_{lk}({\bf H}_{lk,l}{\bf Q}_{lk})^{H}$ is the covariance matrix of the noise plus interference, where ${\bf Q}_{ik}$ is the beamforming matrix at Uik, and ${\bf H}_{ik,l}$ is the uplink channel between Uik and BS $l$. As it can be easily verified, ${\bf S}_{l}$ and ${\bf D}_{l}$ in (50) are quadratic in channels and beamforming matrices, which follow a similar structure as the matrices ${\bf S}_{lk}$ and ${\bf D}_{lk}$ in (13). Thus, to obtain a quadratic and concave surrogate function for $r_{l}$, we can employ the bounds in Lemma 3 and Lemma 4, and follow the steps in the proof of Lemma 5. Once the surrogate functions for the rates are calculated, the beamforming matrices and RIS coefficients can be updated by solving the corresponding optimization problems according to our proposed framework.
@@ IV Numerical results
In this section, we provide numerical results based on Monte Carlo simulations. To this end, we consider a two-cell system with one RIS and $K$ users per cell, similar to [36, Fig. 2]. We also consider that each BS/user/RIS has $N_{BS}/N_{u}/N_{RIS}$ TAs/RAs/elements. Moreover, the locations and heights of the users/BSs/RISs are chosen similar to [36]. We assume that the power budgets of the BSs are equal to $P$. To generate the channels, we assume that the links with respect to the RISs benefit from a line of sight (LoS) for both BSs and the users that are located in the same cell as the RIS. Therefore, these links follow Rician fading associated with a Rician factor of $3$. More particularly, these channels are generated according to [28, (60)-(60)]. On the other hand, we assume that the direct links between the BSs and the users as well as the links through the RISs across the cells are of a non-LoS (NLoS) nature and consequently, follow a Rayleigh distribution. This means that each entry of the corresponding channel matrices follows a complex-valued proper Gaussian distribution with zero mean and unit variance. The large-scale fading of the links is modeled according to [28, (59)]. The path-loss exponents of the LoS and NLoS links are $2.2$ and $3.75$, respectively. The noise power density and channel bandwidth are, respectively, assumed to be $-174$ dBm per Hz and $1.5$ MHz. The other simulation parameters are chosen according to [28, 36].
In the following, we provide numerical results for SE and EE maximization in Section IV-A and Section IV-B, respectively. The schemes considered in this section are as follows:
  • RIS (or RISI): Our algorithms for MIMO RIS-assisted URLLC systems with multiple data streams per users, and $\mathcal{T}_{U}$ (or $\mathcal{T}_{I}$).
@@
As emphasized in Section I, there is no other work on multi-user MIMO RIS-aided systems with FBL coding. Thus, we compare the performance of our proposed algorithm to a single-stream data transmission scheme, the multiple-stream data transmission scheme for systems without RIS, and with a non-optimized RIS coefficients as benchmarks.
@@ IV-A1 Impact of power budget
In Fig. 2, we show the average max-min rate and sum rate versus $P$ for $N_{{BS}}=4$, $N_{u}=4$, $K=2$, $L=2$, $M=2$, $n_{t}=256$, $\epsilon=10^{-5}$, and $N_{{RIS}}=20$. In this figure, the RISs significantly improve the min-max rate and the sum rate, when the RIS elements are optimized. Surprisingly, employing RISs having random elements degrades the max-min rate. Moreover, we can observe that the benefits of RIS increase with $P$ for these examples. Additionally, the multi-stream scheme substantially improves the SE for these two examples, where even the No-RIS scheme is slightly better than the single-stream scheme, especially at high SNRs. In these examples, the maximum number of streams per user is $I=\min(N_{{BS}},N_{u})=4$, which makes single-stream transmission gravely suboptimal. Indeed, when the system complexity increases, a more sophisticated resource allocation scheme should be adopted in order to avoid performance degradation.
@@ IV-A2 Impact of packet length
Fig. 3 shows the average max-min rate and the relative performance improvement by RIS versus $n_{t}$ for $P=10$ dB, $N_{{BS}}=4$, $N_{u}=3$, $K=2$, $L=2$, $M=2$, $\epsilon=10^{-5}$, and $N_{{RIS}}=20$. Observe that an RIS can substantially increase the average max-min rate. The benefits of deploying RISs decrease with $n_{t}$. Indeed, the shorter the packet length is, the higher the relative improvements provided by RISs can be. As indicated, the packet length correlates with the level of stringency in the latency constraint. Shorter packet lengths are needed, when the latency constraint is more stringent. Thus, this result shows that the RIS benefits increase as the latency constraint becomes more stringent. In other words, RISs can even be more beneficial for URLLC systems. Moreover, the rates increase with $n_{t}$ and converge to the Shannon rate when $n_{t}$ becomes higher. Furthermore, we can observe that employing multi-stream data transmission substantially increases the average max-min rate for all the values of $n_{t}$. Note that the benefits of multi-stream data transmission in this figure is lower than in Fig. 2 since the maximum number of streams per users, $I$, is $3$ in this example, which is less than $I$ in Fig. 2, and the benefits of multi-stream schemes grows with $I$.
@@ IV-A4 Comparison of reflective RIS and STAR-RIS
IV-A5 Convergence behavior
Figure 8: Average max-min rate versus the number of iterations for $N_{BS}=4$, $N_{u}=3$, $K=2$, $\epsilon=10^{-4}$, $n_{t}=256$ bits and $N_{RIS}=20$.
Fig. 8 shows the average max-min rate versus the number of iterations for $N_{BS}=4$, $N_{u}=3$, $K=2$, $\epsilon=10^{-4}$, $n_{t}=256$ bits and $N_{RIS}=20$. This figure illustrates the convergence of the considered algorithms, from which the performance-complexity tradeoff may be inferred. In this example, our scheme proposed for RIS-aided systems employing multiple streams per user outperforms the final solution of the other algorithms after as few as four iterations. Indeed, the lower computational complexity of these alternative algorithms results in significant performance degradation, while to achieve a certain target performance, the RIS scheme requires substantially fewer iterations. Additionally, the No-RIS, RIS-Rand, and SS-RIS algorithms may not support a specific max-min rate, while our framework can achieve it with a reasonable number of iterations. For instance, while the No-RIS, RIS-Rand, and SS-RIS algorithms struggle to support a max-min rate of 8 b/s/Hz, our framework achieves it in only nine iterations.
@@ IV-B1 Impact of $P_{c}$
Fig. 9 shows the average max-min EE and GEE versus $P_{c}$ for $N_{{BS}}=5$, $N_{u}=5$, $K=2$, $L=2$, $M=2$, $n_{t}=256$ bits, $\epsilon=10^{-5}$, and $N_{{RIS}}=20$. Note that we use the term max-min EE to refer to the highest minimum EE. As it can be observed, RISs significantly increase the average max-min EE and GEE with FBL. Interestingly, RISs may reduce the EE, if their elements are random. However, our proposed algorithms can statistically enhance the EE of RIS-aided scenarios. For instance, in the particular example of Fig. 9a, an RIS provides more than $10\%$ improvements over the systems disregarding the RISs for all the values of $P_{c}$ considered. Additionally, the multi-stream scheme significantly outperforms the single-stream data transmission in the both examples.